OpenStudy (anonymous):

Write an equation for the nth term of the given arithmetic sequence. -13, -28, -43, -59,

OpenStudy (tkhunny):

Oops, it's not an arithmetic squence. Please prove it.

OpenStudy (anonymous):

Agreed, I don't see the sequence

OpenStudy (anonymous):

well thats all the question told me but the answers they gave me are a) a^n = -15n - 28 b) a^n = -13n + 1 c) a^n = -15n + 2 d)a^n = 15n - 28

OpenStudy (anonymous):

Just use brute force. Start plugging n's in each choice and see which one results in a matching sequence

OpenStudy (tkhunny):

Don't try to solve it. The answer is "None of the Above". It is NOT an arithmetic sequence.

OpenStudy (anonymous):

Ok thanks!

OpenStudy (tkhunny):

Do you know why this is so?

OpenStudy (anonymous):

Actually, the brute force method is incorrect also

OpenStudy (tkhunny):

-13, -28, -43, -59, -28 - (-13) = -15 <== This could be the common difference. -43 - (-28) = -15 <== Okay, I almost believe it. -59 - (-43) = -16 <== That's no good. We do NOT have a common difference. If it has said -58, we would be in business.

OpenStudy (anonymous):

Exactly, what I was about to ask. Should that be 58?

OpenStudy (anonymous):

-58 I meant

OpenStudy (anonymous):

It might be! they might have had a typo or something. ill try it with -58 and see if i get an answer

OpenStudy (tkhunny):

If it were me on an exam, I would first state that it is not an arithmetic sequence - and prove it. I would then change the last value to -58 and work the problem as if it were so intended.